Fermi National Laboratory
 

Operation Summary for May 2, 2011 - May 4, 2011

 

Monday May 2

The day shift began with Operations monitoring Tevatron (TEV) store 8707 (193.22E30) and the Recycler stash (148.8E10), with the Antiproton Source (Pbar) stacking (11.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported that the MI beam was briefly interrupted when an interlocked door was inadvertently opened. After operators reset the MI interlocks, NuMI resumed running with the horn on. MI and RF personnel conducted coalescing checks.

During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations terminated store 8707 at 11:11 PM and put the TEV into shot setup. Operations halted stacking so that controls personnel could work on a Pbar repeater. Operations reported some short downtimes due to Linac and Booster RF stations (LRF1 & 2 and BRF12 & 18).

Tuesday May 3

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (220.0E10), with Pbar stacking halted for repeater work (8.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, controls personnel reported that the Pbar CAMAC link problem was due to a loose connector. Operations established store 8708 at 12:35 AM with an average initial luminosity (AIL) of 401.9E30. Operations established beam to T-978 at 4:17 AM.

During the day shift, Operations monitored the accelerator complex.

During the evening shift, Operations reported that the TEV sector F0 LCW air compressor tripped off at 5:53 PM. FESS personnel reset the compressor. Operations disabled beam to T-978 at 5:59 PM. Operations terminated store 8708, put the TEV into shot setup, and established store 8709 at 9:36 PM with and AIL of 431.07E30. This store set a New TEV Record for initial luminosity. At 9:46 PM, Operations reported that an emittance blowup for protons and antiprotons in the TEV caused a loss of 50E30 from the store. At 10:03 PM, Operations reported losing a transfer of antiprotons due to a Recycler kicker (R:KPS2A) prefire indication. Operations reported short downtimes during the shift due to problems with LRF5, a Klystron RF station (KRF1), BRF14, and BRF18.

Wednesday May 4

The midnight shift began with Operations monitoring store 8709 (223.62E30) and the stash (72.9E10), with Pbar stacking (21.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 2:39 AM, Operations reported that bad beam positions at the MI-8 collimators were causing MiniBooNE trips during antiproton transfers to the Recycler. Operations established beam to T-978 at 4:36 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TEV personnel to conduct parasitic studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/04/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory